OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations eliminated today? A)1300 B)1120 C)708 D)412

OpenStudy (anonymous):

Come on anyone?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

No A

OpenStudy (anonymous):

Oh wait could be B

OpenStudy (anonymous):

D looks viable...

OpenStudy (anonymous):

Thanks...

OpenStudy (phi):

this looks like algebra. start with number of one-step equations = x (give the number of 1-step questions a "name") number of two-step equations = y It says the # of 1 step plus # of 2 step add up to equal to 1120. you would say x+y= 1120

OpenStudy (phi):

now translate If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, into math: three times the number of one-step equations 3 * x or just 3x minus means - so you now have 3x - twice the number of two-step 2 * y or 2y you now have: 3x - 2y is equal to 1300, = 1300 3x-2y = 1300

OpenStudy (phi):

you have 2 equations and 2 unknowns x+y= 1120 3x-2y = 1300 how many two-step equations eliminated today? they want you to solve for y

OpenStudy (anonymous):

So those are the two equations i have to do to find y?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thank you so much!